# NoTeeline: Supporting Real-Time Notetaking from Keypoints with Large Language Models

### v1

@@ Acknowledgements
We thank Dan Saffer and Chris Maury for their valuable feedback on the design of NoTeeline, which contributed significantly to its development. We also appreciate the insightful comments and suggestions provided by Pranav Khapde, Nazmus Saquib, and Yi-Hao Peng on earlier drafts of this work.
@@ Appendix A Participants’ Information
Table 3. Demography of user study participants.
Group No. # of Participants Tool_1 Video_1 Tool_2 Video_2
@@ B.1. Video Clip Selection for Onboarding Session
We carefully selected three videos for the onboarding from three different domains: Science, Psychology, and blogs. We extracted 15-20 seconds snippet from each video and stored the timestamps. The videos are listed below: 1) How Mitochondria Produce Energy (Science); 2) How to get motivated even when you don’t feel like it (Psychology); 3) Apple Vision Pro VS Meta Quest 3! HONEST Comparison! (Blog).
@@ B.2. Video Selection Criteria for User Study
We picked two videos for the two conditions. The videos met the following criteria: 1) not too long ($<5$ minutes); 2) information heavy where users are exposed to scientific concepts (– so that they feel the need of taking note); 3) interesting to a wide range of audience. Based on these criteria, we selected two TED-Ed videos entitled ‘How memories form and how we lose them’ (url1) and ‘Is it possible to lose weight fast?’ (url2). Both videos relate to Human Biology, are ~4 minutes long, and have been viewed over 2 million times, reflecting their wide appeal.
@@ Appendix E Questionnaire for NeedFinding Study
  • What is your typical notetaking process when watching online informational videos (e.g., podcasts, lecture, tutorials)?
@@
  • If there is an AI assisted system that could take your manual notes and organize them automatically for you, would you like to use them?
@@
4. Just write a single note point, users will request repeatedly for new points they want to add.
@@
Make sure that the note aligns with the user’s writing style, so that they can read it easily. Use the same writing style as shown below. Here are three examples:
@@
[] Given a list of points, Your task is to perform topic modeling over them. Arrange the points into topics and provide a name to each individual topic. Please mark the topic within <Topic></Topic>tags. the points are marked with <p>tag as well. Follow these rules:
1. Each topic should have more than one points.
@@
Given a topic description, your task is to generate five multichoice question with answer. Please mark the question within <Question></Question>tags, individual choices within <Choice></Choice/>tags and answer within <Answer></Answer>tags. Make sure not to put the write choice in the same choice option, randomly assign it within A, B, C or D. Here is an example:
Topic: Resilience refers to how well you can deal with and bounce back from the difficulties of life. It can mean the difference between handling pressure losing your cool. Resilient people tend to maintain a more positive outlook and cope with stress .. (cont.)..
@@
Hallucination Rate 4.76 5.88 11.11 6.67 6.67 9.09 0.0 0.0 10.71 13.33 10.0 3.33
Factual Consistency Rate 95.24 94.12 88.89 93.33 93.33 90.91 100.0 100.0 89.29 86.67 90.0 96.67
Table 7. Correctness of generated note for each participant in terms on Hallucination Rate (%, lower is better) and Factual Consistency Rate (%, higher is better)

### v2

@@ Acknowledgements
We thank Dan Saffer and Chris Maury for their valuable feedback on the design of NoTeeline, which contributed significantly to its development. We also appreciate the insightful comments and suggestions provided by Pranav Khapde, Nazmus Saquib, Riku Arakawa, Yi-Hao Peng and Yaxin Hu on earlier drafts of this work.
@@ Appendix A Participants’ Information
Figure 12. Demography of user study participants.
Group No. # of Participants Tool #1 Video #1 Tool #2 Video #2
@@ B.1. Video Clip Selection for Onboarding Session
We carefully selected three videos for the onboarding from three different domains: Science, Psychology, and blogs. We extracted 15-20 second snippets from each video and stored the timestamps. The videos are listed below: 1) How Mitochondria Produce Energy (Science); 2) How to get motivated even when you don’t feel like it (Psychology); 3) Apple Vision Pro VS Meta Quest 3! HONEST Comparison! (Blog).
@@ B.2. Video Selection Criteria for User Study
We picked two videos for the two conditions. The videos met the following criteria: 1) not too long ($<5$ minutes); 2) information heavy where users are exposed to scientific concepts (– so that they feel the need to take note); 3) interesting to a wide range of audience. Based on these criteria, we selected two TED-Ed videos entitled ‘How memories form and how we lose them’ (url1) and ‘Is it possible to lose weight fast?’ (url2). Both videos relate to Human Biology, are ~4 minutes long, and have been viewed over 2 million times, reflecting their wide appeal.
@@ Appendix E Questionnaire for NeedFinding Study
  • What is your typical notetaking process when watching online informational videos (e.g., podcasts, lectures, tutorials)?
@@
  • If there is an AI-assisted system that could take your manual notes and organize them automatically for you, would you like to use them?
@@
4. Just write a single note point, and users will request repeatedly for new points they want to add.
@@
Make sure that the note aligns with the user’s writing style so that they can read it easily. Use the same writing style as shown below. Here are three examples:
@@
[] Given a list of points, Your task is to perform topic modeling over them. Arrange the points into topics and provide a name for each topic. Please mark the topic within <Topic></Topic>tags. the points are marked with <p>tag as well. Follow these rules:
1. Each topic should have more than one point.
@@
Given a topic description, your task is to generate five multichoice questions with answers. Please mark the question within <Question></Question>tags, individual choices within <Choice></Choice/>tags and answer within <Answer></Answer>tags. Make sure not to put the right choice in the same choice option, randomly assign it within A, B, C or D. Here is an example:
Topic: Resilience refers to how well you can deal with and bounce back from the difficulties of life. It can mean the difference between handling pressure and losing your cool. Resilient people tend to maintain a more positive outlook and cope with stress .. (cont.)..
@@
Hallucination Rate (%) 4.76 5.88 11.11 6.67 6.67 9.09 0.0 0.0 10.71 13.33 10.0 3.33
Factual Consistency Rate (%) 95.24 94.12 88.89 93.33 93.33 90.91 100.0 100.0 89.29 86.67 90.0 96.67
Table 6. Correctness of generated note for each participant in terms of Hallucination Rate (%, lower is better) and Factual Consistency Rate (%, higher is better)